How do I solve for N in the following equation?
$45 = [N($50) + $25]/(N + 1)
Thank you.

Found 2 solutions by josgarithmetic, richwmiller:
Answer by josgarithmetic(39617)
Answer by richwmiller(17219)

45 = (N(50) + 25)/(N + 1)
45 = (50N+25)/(N + 1)
45(N+1)=50N+25
45N+45=50N+25
20=5N
4=N